OpenStudy (anonymous):

What is the percent of increase from 50 to 90? A. 40% B. 44% C. 56% D. 80%

OpenStudy (anonymous):

(90-50)/90=40/90 for %=(40/90)*100=400/9=44.44

OpenStudy (anonymous):

roundin give 44%

OpenStudy (anonymous):

% change = (100%)(new-old)/old

OpenStudy (anonymous):

What is the percent of decrease from 200 to 110? A. 45% B. 55% C. 82% D. 90%

OpenStudy (anonymous):

Hold it. % increase from 50 to 90 is not 44% (90-50)/50= 80%

OpenStudy (anonymous):

The price of a suit decreased from $280 to $182 during a clearance sale. What is the percent of decrease in the price? Round to the nearest percent, if necessary. A. 65% B. 35% C. 54% D. 98%